a) The correct precedence relationship with one of the possible assignment of tasks to workstations is shown in ?
b) What is the operating efficiency of the assembly​ line, using your answer to​ (a)?
c) What is the theoretical minimum number of​ workstations?
d) What is the idle time per boat​ produced?
IOIIGWufk. Cny Score: 0 of 1 pt Problem 9.18 Assembly Line Balancing 10 of 11 (6 complete) HW Score: 54.55%, 6 of 11 Question Holp The final assembly plant is in Cupertino, California. The assembly area is available for Fig. 1 production of the Mach 10 for 200 minutes per day. (The rest of the time it is busy making other products.) The daily demand is 64 boats. Given the following Time (mins) Predecessors Fig. 2 D, E, F G, H a) The correct precedence relationship with one of the possible assignment of tasks to workstations is shown in Fig. 3
media%2F802%2F80243684-66c8-4ca2-ad01-68
media%2Fc5d%2Fc5d76152-c99c-4a54-87b9-f3
Answer #1

a. Both Fig 2 and 3 follow the precedence relationship. How ever, fig 3 only qualified for the workstation assignment based on the cycle time.

b. Efficiency= total of task times/(number of work station* cycle time)= 14/(3*5)= 93%.

c. Theoretical cycle time= available time/ output= 300/58= 5.11

Number of work station= Total of task times/ cycle time= 14/5.11= 2.73= 3 work stations.

d. Idle time is on work station 3= 1 minute.

e. For 400 minutes available, Cycle time= 400/58= 6.89.

Number of work stations= total of task times/ cycle time= 14/6.89= 2.03= 3 work stations.

f. Efficiency for this case= Total of task times/ (Number of work stations* cycle time)= 14/(3*6.89)= 66.67%.


answered by: ANURANJAN SARSAM